OpenStudy (peterg2001):

The measures 9, 12, and 15 create a right triangle. True False

OpenStudy (peterg2001):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

use pythgorian theorem: 9^2+12^2=15^2 this is obviously false

OpenStudy (welshfella):

15^2 = 9^2 + 12^2 225 = 81 + 144 225 = 225 its TRUE